# Quadratic-form Optimal Transport

### v2

@@ 4.1 Type-XY product of two quadratic cost functions
$c(x,y,x^{\prime},y^{\prime})=f(x,y)g(x^{\prime},y^{\prime}),\mbox{ \leavevmode\nobreak\ \leavevmode\nobreak\ where $f,g$ are quadratic functions},$ (4.1)
@@
Since the quadratic function $z\mapsto 8z^{2}-2z-10$ is minimized at $z=1/8$, we see that if $\mathrm{Cov}(X,Y)=1/8$, the law $\pi$ of $(X,Y)$ is a QOT minimizer. This is achieved, for example, by $\pi=(9/16)\pi_{\mathrm{com}}+(7/16)\pi_{\mathrm{ant}}$, where $(X,X)\buildrel\mathrm{law}\over{\sim}\pi_{\mathrm{com}}$ and $(X,-X)\buildrel\mathrm{law}\over{\sim}\pi_{\mathrm{ant}}$. On the other hand, since the range of $\mathbb{E}[XY]$ is $[-1,1]$, the unique QOT maximizer is given by the antimonotone coupling $X=-Y$, where $\mathbb{E}[XY]=-1$.
Following the same idea, the next result replaces the quadratic functions $f,g$ in Proposition 4.1 by submodular functions $f,g$ that are identical up to a constant term.
@@
$c(x,y,x^{\prime},y^{\prime})=(f(x,y)+a_{1})(f(x^{\prime},y^{\prime})+a_{2}),\mbox{ \leavevmode\nobreak\ \leavevmode\nobreak\ where $f$ is submodular and $a_{1},a_{2}\in\mathbb{R}$},$ (4.2)
@@
The same conclusion of Proposition 4.3 holds true if we consider the cost function $c(x,y,x^{\prime},y^{\prime})=(f(x,y)+a_{1})(-f(x^{\prime},y^{\prime})+a_{2})$, or supermodular $f$ instead of submodular $f$. We omit these simple variants.
@@ 4.2 Jointly submodular cost functions
    Generalizing (i), suppose that $c(x,y,x^{\prime},y^{\prime})=h(c_{1}(x,y),c_{2}(x^{\prime},y^{\prime}))$ and $c\in\mathcal{C}(\mu,\nu)$, where both $c_{1},c_{2}$ are submodular and componentwise increasing, and $h:\mathbb{R}^{2}\to\mathbb{R}$ is componentwise increasing and concave. We can check that the submodularity condition in Theorem 4.4 is satisfied. Hence, the comonotone coupling is a minimizer. Similarly, the antimonotone coupling is a maximizer if $-c\in\mathcal{C}(\mu,\nu)$. This includes, for instance, $c(x,y,x^{\prime},y^{\prime})=\min\{c_{1}(x,y),c_{2}(x^{\prime},y^{\prime})\}$, and $c(x,y,x^{\prime},y^{\prime})=(c_{1}(x,y)+c_{2}(x^{\prime},y^{\prime}))^{p}$ where $p\in(0,1)$ and $c_{1},c_{2}$ nonnegative.
@@
    The function $|x+y-x^{\prime}-y^{\prime}|$ is related to the Gini coefficient (see Example 6.5 below) and satisfies the supermodularity condition in Theorem 4.4, since $(x,y)\mapsto|x+y+c|$ is supermodular for every $c\in\mathbb{R}$. More generally, the cost function $|x+y-x^{\prime}-y^{\prime}|^{p}$ (resp. $|x-y+x^{\prime}-y^{\prime}|^{p}$) for $p\geqslant 1$ induces the antimonotone (resp. comonotone) coupling as an optimizer. Similarly, the antimonotone coupling is an optimizer of QOT with cost function $|x+y+x^{\prime}+y^{\prime}|^{p}$ for $p\geqslant 1$.
@@
    The cost function $c(x,y,x^{\prime},y^{\prime})=\mathrm{sgn}(x-x^{\prime})\,\mathrm{sgn}(y-y^{\prime})$ defines Kendall’s tau; see Example 6.4 below. It is elementary to verify that $(x^{\prime},y^{\prime})\mapsto c(x,y,x^{\prime},y^{\prime})$ and $(x,y)\mapsto c(x,y,x^{\prime},y^{\prime})$ are supermodular, and hence the transport cost is maximized by the comonotone coupling and minimized by the antimonotone coupling.
@@ 4.3 Gromov–Wasserstein-type cost functions
We now consider a family of type-XX cost functions, called the GW-type cost functions (see Example 6.6 in Section 6 for the GW distance). In what follows, we say that $\nu$ is an increasing (resp. decreasing) location-scale transform of $\mu$ if $\nu=\mu\circ\ell^{-1}$ for some strictly increasing (resp. decreasing) linear map $\ell:\mathbb{R}\to\mathbb{R}$; that is, $\ell(x)=ax+b$ for some $a>0$ (resp. $a<0$) and $b\in\mathbb{R}$.
@@
    The choice $h(u,v)=|u^{q}-v^{q}|^{p}$ corresponds to the $(p,q)$-GW transport cost, see (6.3). One can verify that $h$ is submodular if $p\geqslant 1,q>0$ on $[0,\infty)^{2}$. Theorem 4.6 then implies that in the case $\mu=\nu$, $\pi_{\mathrm{com}}$ (and $\pi_{\mathrm{ant}}$ if $\mu$ is symmetric) is a minimizer for (6.3). This aligns with the intuition that the GW distance measures distances between metric measure spaces.
@@ 4.4 A special class of separable cost functions and the V-transport
For instance, if $\mu,\nu\buildrel\mathrm{law}\over{\sim}\mathrm{U}$, then $\pi_{\rm v}$ is the distribution of $(U,|2U-1|)$ where $U\buildrel\mathrm{law}\over{\sim}\mathrm{U}$; see Figure 1(b). If $\nu$ is atomless with median $m_{\nu}$, then $\pi_{\rm v}$ is the arithmetic average of the antimonotone coupling of $\mu$ and $2\nu|_{(-\infty,m_{\nu}]}$, and the comonotone coupling of $\mu$ and $2\nu|_{(m_{\nu},\infty)}$.
@@ 5.1 The rectangular cost function
For $(x,y),(x^{\prime},y^{\prime})\in\mathbb{R}^{2}$, denote by $[(x,y),(x^{\prime},y^{\prime})]$ the unique closed rectangle in $\mathbb{R}^{2}$ whose sides are parallel to the $xy$-axes and two of whose corners are given by $(x,y),(x^{\prime},y^{\prime})$ if $x\neq x^{\prime}$ and $y\neq y^{\prime}$, and the empty set otherwise. It holds that for $\pi\in\Pi(\mu,\nu)$,
@@
Our goal is to show that the integrand in (5.3) is uniquely minimized for all $(u,v)$ by $\pi_{\mathrm{dia}}=\pi_{\mathrm{dia}}$, which suffices for our purpose: since $\mu,\nu\in\mathcal{P}_{1}(\mathbb{R})$, the transport cost for the independent coupling is finite, and hence so is the transport cost for $\pi_{\mathrm{dia}}$. For a fixed $(u,v)\in\mathbb{R}^{2}$ such that $\mu$ has no atom at $u$ and $\nu$ has no atom at $v$, define $A(u,v)=\pi((-\infty,u]\times(-\infty,v])$. Also, denote by $F_{\mu},F_{\nu}$ the distribution functions of $\mu,\nu$, so $F_{\mu}$ is continuous at $u$ and $F_{\nu}$ is continuous at $v$. Using $\pi\in\Pi(\mu,\nu)$, we have
@@
The absolute value in the cost function $c$ in Theorem 5.2 is important. If $c$ is specified by $c(x,y,x^{\prime},y^{\prime})=(x-x^{\prime})(y-y^{\prime})$ (or equivalently, if the objective is $\mathrm{Cov}(X,Y)$; c.f. Example 6.3 below), then the minimizer is the antimonotone coupling instead of the diamond transport.
@@ 5.2 A class of type-XX cost functions with convex QOT
for $\alpha\in(0,1/2]$. Specifically, one of our assumptions in this result is that the quadratic-form transport cost is convex in the transport plan $\pi$. Schoenberg’s theory of complete monotonicity provides a convenient sufficient condition for such convexity. A nonnegative continuous function ϕ:++:italic-$\phi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is completely monotone if $\phi$ is $C^{\infty}$ on $(0,\infty)$ and satisfies $(-1)^{n}\phi^{(n)}(u)\geqslant 0$ for $n\geqslant 0,\,u>0$ (see (Berg et al., 1984, Section 4.6)). In particular, $\phi$ is bounded and decreasing. We use the standard calculus notation $\phi^{\prime}$ for the first derivative $\phi^{(1)}$, which should not be confused with the apostrophe in $x^{\prime},y^{\prime}$.
@@
for some completely monotone function ϕ:+:italic-$\phi:\mathbb{R}_{+}\to\mathbb{R}$ satisfying
@@
[…]
@@
    Suppose that $\mu,\nu$ are supported on $[-1/2,1/2]$. Let $\phi(u)=(\beta+u)^{-\gamma}$, where $\gamma>0$ and $\beta>2\gamma+1$. This leads to the cost function
@@
    Suppose that $\mu,\nu$ are supported on $[-L,L]$ where $L>0$. Let $p\in(1/2,1)$ and $\phi(u)=e^{-\alpha u^{p}}$, where $0<\alpha\leqslant(2p-1)/(2p(4L^{2})^{p})$. The cost function is then given by
@@
We first claim that the resulting QOT problem is convex in $\pi$. Since $\phi$ is completely monotone, the function $\psi(u):=\phi(\sqrt{u})$ is a continuous positive definite function on $\mathbb{R}_{+}$ by Schoenberg’s theorem (Theorem 4.6.13 and Example 5.1.3 of Berg et al. (1984)). In this way, we write $c(x,y,x^{\prime},y^{\prime})=\psi(|x-x^{\prime}|)\psi(|y-y^{\prime}|)$ where $\psi$ is positive definite. By the Schur product theorem, the product of two positive definite functions is also positive definite (see Theorem 3.1.12 of Berg et al. (1984)). Therefore, $c(x,y,x^{\prime},y^{\prime})$ is a positive definite kernel in the two variables $(x,y),(x^{\prime},y^{\prime})\in\mathbb{R}^{2}$. Proposition 3.1 then implies the QOT problem is convex in $\pi$.
Since the cost function (5.4) is translation-invariant, we may assume that $\mu,\nu$ are both symmetric along 00. We next verify
@@
Consider the classic OT problem with cost function $\tilde{c}$ given by (5.7). Denote a minimizer by $\pi_{\tilde{c}}$. By linearity of the classic OT problem and (5.8), another minimizer is given by the symmetrized version $\hat{\pi}_{\tilde{c}}$ of $\pi_{\tilde{c}}$: the law of the uniform mixture of $(X,Y),\,(-X,Y),\,(X,-Y),$ and $(-X,-Y)$ where $(X,Y)\buildrel\mathrm{law}\over{\sim}\pi_{\tilde{c}}$. Therefore, $\hat{\pi}_{\tilde{c}}$ is symmetric along the $x$ and $y$ axes. By cyclical monotonicity and Lemma 5.8, any minimizer is antimonotone on the first and third quadrants and comonotone on the second and fourth quadrants. This implies that $\hat{\pi}_{\tilde{c}}=\pi_{\mathrm{dia}}$, and hence $\pi_{\mathrm{dia}}:=\pi_{\mathrm{dia}}$ is a minimizer of the OT problem with cost function $\tilde{c}$.
@@
The final claim on uniqueness follows from the fact that if $\phi$ is non-constant, the cost function $c$ is strictly positive definite in the two variables $(x,y),(x^{\prime},y^{\prime})\in\mathbb{R}^{2}$ (Theorem 3’ of Schoenberg (1938)), and hence the transport cost is strictly convex in $\pi$ with a unique minimizer. ∎
@@ 5.3 The $q$-rectangular cost function
Fix $q\in(1,2]$ and assume without loss of generality that $\mu,\nu$ are both symmetric around 00. Assume first that $\mu,\nu$ are supported in $[-L,L]$ for some $L>0$. For $\alpha>0$, consider the cost function
@@
uniformly on $[-L,L]^{4}$ as $\alpha\to 0$. Therefore, the QOT problem with cost function $(|x-x^{\prime}|^{q}+|y-y^{\prime}|^{q})^{2}/2$ also has $\pi_{\mathrm{dia}}$ as a minimizer. Removing the QOT-irrelevant terms, we see that this is also the case for the cost function $|(x-x^{\prime})(y-y^{\prime})|^{q}$, as desired.
@@
6 Examples of QOT
In this section, we discuss several examples of QOT that appear in different fields.
Example 6.1 (Sum of bivariate functions).
Example 6.2 (Variance minimization with given marginals).
$\iint c\,\mathrm{d}\pi\otimes\mathrm{d}\pi&=\begin{cases}\mathbb{E}[f(X,Y)^{2}]-\mathbb{E}[f(X,Y)]^{2}&\text{ if }\mathbb{E}[f(X,Y)^{2}]<\infty\\ \infty&\mbox{ otherwise}\end{cases}\\ &=\mathrm{Var}(f(X,Y)).\end{aligned}$ (6.1)
Example 6.3 (Covariance).
For $(X,Y)\buildrel\mathrm{law}\over{\sim}\pi$, we can verify
$=\mathbb{E}[XY]-\mathbb{E}[X]\mathbb{E}[Y]=\mathrm{Cov}(X,Y).$
Example 6.4 (Kendall’s tau).
$\tau=\mathbb{E}[\mathrm{sgn}((X-X^{\prime})(Y-Y^{\prime}))],$
where $(X^{\prime},Y^{\prime})$ is an independent copy of $(X,Y)$. Intuitively, it equals the probability of concordance minus that of discordance between $(X,Y)$ and $(X^{\prime},Y^{\prime})$. Clearly, $\tau$ of $(X,Y)\buildrel\mathrm{law}\over{\sim}\pi$ can be written as the quadratic-form transport cost with the type-XX cost function $c(x,y,x^{\prime},y^{\prime})=\mathrm{sgn}(x-x^{\prime})\,\mathrm{sgn}(y-y^{\prime})$. For given marginals $\mu,\nu$, it is well-known that $\tau(\pi)$ over $\pi\in\Pi(\mu,\nu)$ is maximized by the comonotone coupling with maximum value $1$ and minimized by the antimonotone coupling with minimum value $-1$ (this can also be checked by Theorem 4.4; see Example 4.5). Another equivalent formulation is $\tau(\pi)=4\int\pi\mathrm{d}\pi-1,$ from which the quadratic form in $\pi$ is visible.
Example 6.5 (Gini deviation and Gini coefficient).
where $Z^{\prime}$ is an independent copy of $Z$. The value $\mathrm{GD}(Z)$ is called the Gini deviation of $Z$, and $\mathrm{GC}(Z)$ is called the Gini coefficient of $Z$, both of which are commonly used as measures of distributional variability or inequality in economics and risk management; see e.g., Gastwirth (1971) and Furman et al. (2017). Similarly to the variance in (6.1), minimization of $\mathrm{GD}(f(X,Y))$ for some measurable $f:\mathbb{R}^{2}\to\mathbb{R}$ over $X\buildrel\mathrm{law}\over{\sim}\mu$ and $Y\buildrel\mathrm{law}\over{\sim}\nu$ can be written as the QOT problem with the type-XY cost function $|f(x,y)-f(x^{\prime},y^{\prime})|$. Moreover, the minimization of $\mathrm{GD}(X+Y)$ can be written as
which is equivalent to a QOT problem with cost function $c(x,y,x^{\prime},y^{\prime})=|x+y-x^{\prime}-y^{\prime}|$, noting that the denominator of (6.2) does not involve $\pi$. This cost function is both type-XX and type-XY.
Example 6.6 (Gromov–Wasserstein (GW) distance).
A special case of QOT is the GW distance, a measure of the distance (or similarity) between two metric measure spaces first proposed by Mémoli (2011). Suppose that $(\mathfrak{X},d_{\mathfrak{X}},\mu)$ and $(\mathfrak{Y},d_{\mathfrak{Y}},\nu)$ are metric measure spaces. For $p,q\geqslant 1$, the $(p,q)$-GW distance is defined as
This is an increasing transform of the minimum transport cost of QOT with a type-XX cost function. The GW distance satisfies the triangle inequality and defines a pseudo-metric on metric measure spaces (and a metric on isomorphism classes of metric measure spaces; see Sturm (2023, Corollary 9.3)). The GW distance is a widely used technique in data science, machine learning, computer vision, and computer graphics to align heterogeneous data sets or images (Mémoli (2011); Peyré and Cuturi (2019)). However, in general, solving for the GW distance is a challenging task. In Theorems 5.2 and 5.9, we explicitly characterize the maximizers of the transport cost that appears in (6.3) for $p=2$, $q\in[1,2]$, $\mathfrak{X}=\mathfrak{Y}=\mathbb{R}$ equipped with the Euclidean distance, and symmetric marginals, where the symmetry is not needed for $q=1$.
Example 6.7 (Quadratic Assignment Problem (QAP)).
If the probability measures $\mu,\nu$ are each uniformly distributed on $N$ points, the Monge QOT problem reduces to QAP, which was first introduced by Koopmans and Beckmann (1957) as a model for the allocation problem of indivisible economic activities. Recall the Koopmans–Beckmann problem from (1.5). The work of Lawler (1963) proposed a generalization of the Koopmans–Beckmann QAP, where one solves
$\min_{\sigma\in S_{n}}\sum_{i=1}^{n}\sum_{j=1}^{n}d_{ij\sigma_{i}\sigma_{j}},$
subject to $\pi\in\Pi(\mu,\nu),$

### v4

@@ 5.1 Type-XY product of two quadratic cost functions
[…]
@@
Since the quadratic function $z\mapsto 8z^{2}-2z-10$ is minimized at $z=1/8$, we see that if $\mathrm{Cov}(X,Y)=1/8$, the law $\pi$ of $(X,Y)$ is a QOT minimizer. This is achieved, for example, by $\pi=(9/16)\pi_{\mathrm{com}}+(7/16)\pi_{\mathrm{ant}}$, where $(X,X)\buildrel\mathrm{law}\over{\sim}\pi_{\mathrm{com}}$ and $(X,-X)\buildrel\mathrm{law}\over{\sim}\pi_{\mathrm{ant}}$. On the other hand, since the range of $\mathbb{E}[XY]$ is [1,1][-1,1], the unique QOT maximizer is given by the antimonotone coupling X=YX=-Y, where $\mathbb{E}[XY]=-1$.
Proposition 5.1 follows from the following more general result, which replaces the quadratic functions $f,g$ in Proposition 5.1 by submodular functions $f,g$ that are identical up to a constant term.
@@
Suppose that $\mu,\nu\in\mathcal{P}(\mathbb{R})$ and $f$ is submodular. If the cost function $c$ is given by
[…]
@@ 5.2 Jointly submodular cost functions
    Generalizing (i), suppose that $c(x,y,x^{\prime},y^{\prime})=h(c_{1}(x,y),c_{2}(x^{\prime},y^{\prime}))$ and $c\in\mathcal{C}(\mu,\nu)$, where both $c_{1},c_{2}$ are submodular and componentwise increasing, and $h:\mathbb{R}^{2}\to\mathbb{R}$ is componentwise increasing and concave. We can check that the submodularity condition in Theorem 5.4 is satisfied. Hence, the comonotone coupling is a minimizer. Similarly, the antimonotone coupling is a maximizer if $c\in\mathcal{C}(\mu,\nu)$. This includes, for instance, $c(x,y,x^{\prime},y^{\prime})=\min\{c_{1}(x,y),c_{2}(x^{\prime},y^{\prime})\}$, and $c(x,y,x^{\prime},y^{\prime})=(c_{1}(x,y)+c_{2}(x^{\prime},y^{\prime}))^{p}$ where $p\in(0,1)$ and $c_{1},c_{2}$ nonnegative.
@@
    The function $|x+y-x^{\prime}-y^{\prime}|$ is related to the Gini coefficient (see Example 3.5) and satisfies the supermodularity condition in Theorem 5.4, since $(x,y)\mapsto|x+y+c|$ is supermodular for every $c\in\mathbb{R}$. More generally, the cost function $|x+y-x^{\prime}-y^{\prime}|^{p}$ (resp. $|x-y+x^{\prime}-y^{\prime}|^{p}$) for $p\geqslant 1$ induces the antimonotone (resp. comonotone) coupling as an optimizer. Similarly, the antimonotone coupling is an optimizer of QOT with cost function $|x+y+x^{\prime}+y^{\prime}|^{p}$ for $p\geqslant 1$.
@@
    The cost function $c(x,y,x^{\prime},y^{\prime})=\mathrm{sgn}(x-x^{\prime})\,\mathrm{sgn}(y-y^{\prime})$ defines Kendall’s tau; see Example 3.4. It is elementary to verify that $(x^{\prime},y^{\prime})\mapsto c(x,y,x^{\prime},y^{\prime})$ and $(x,y)\mapsto c(x,y,x^{\prime},y^{\prime})$ are supermodular, and hence the transport cost is maximized by the comonotone coupling and minimized by the antimonotone coupling.
@@ 5.3 Gromov–Wasserstein-type cost functions
We now consider a family of type-XX cost functions, called the GW-type cost functions (see Example 3.6 for the GW distance). In what follows, we say that $\nu$ is an increasing (resp. decreasing) location-scale transform of $\mu$ if $\nu=\mu\circ\ell^{-1}$ for some strictly increasing (resp. decreasing) linear map $\ell:\mathbb{R}\to\mathbb{R}$; that is, $\ell(x)=ax+b$ for some $a>0$ (resp. $a<0$) and $b\in\mathbb{R}$.
@@
    The choice $h(u,v)=|u^{q}-v^{q}|^{p}$ corresponds to the $(p,q)$-GW transport cost, see (3.3). One can verify that $h$ is submodular if $p\geqslant 1,q>0$ on $0,\infty)^{2}$. Theorem 5.6 then implies that in the case $\mu=\nu$, $\pi_{\mathrm{com}}$ (and $\pi_{\mathrm{ant}}$ if $\mu$ is symmetric) is a minimizer for (3.3). This aligns with the intuition that the GW distance measures distances between metric measure spaces.
@@ 5.4 A special class of separable cost functions and the V-transport
For instance, if $\mu,\nu\buildrel\mathrm{law}\over{\sim}\mathrm{U}$, then $\pi_{\rm v}$ is the distribution of (U,|2U1|)(U,|2U-1|) where $U\buildrel\mathrm{law}\over{\sim}\mathrm{U}$; see Figure 1(b). If $\nu$ is atomless with median $m_{\nu}$, then $\pi_{\rm v}$ is the arithmetic average of the antimonotone coupling of $\mu$ and $2\nu|_{(-\infty,m_{\nu}]}$, and the comonotone coupling of $\mu$ and $2\nu|_{(m_{\nu},\infty)}$.
@@ 6.1 The rectangular cost function
For $(x,y),(x^{\prime},y^{\prime})\in\mathbb{R}^{2}$, denote by $(x,y),(x^{\prime},y^{\prime})]$ the unique closed rectangle in $\mathbb{R}^{2}$ whose sides are parallel to the $xy$-axes and two of whose corners are given by $(x,y),(x^{\prime},y^{\prime})$ if $x\neq x^{\prime}$ and $y\neq y^{\prime}$, and the empty set otherwise. It holds that for $\pi\in\Pi(\mu,\nu)$,
@@
Our goal is to show that the integrand in (6.3) is uniquely minimized for all $(u,v)$ by ${\pi}=\pi_{\mathrm{dia}}$, which suffices for our purpose: since $\mu,\nu\in\mathcal{P}_{1}(\mathbb{R})$, the transport cost for the independent coupling is finite, and hence so is the transport cost for $\pi_{\mathrm{dia}}$. For a fixed $(u,v)\in\mathbb{R}^{2}$ such that $\mu$ has no atom at $u$ and $\nu$ has no atom at $v$, define $A(u,v)=\pi((-\infty,u]\times(-\infty,v])$. Also, denote by $F_{\mu},F_{\nu}$ the distribution functions of $\mu,\nu$, so $F_{\mu}$ is continuous at $u$ and $F_{\nu}$ is continuous at $v$. Using $\pi\in\Pi(\mu,\nu)$, we have
@@
The absolute value in the cost function $c$ in Theorem 6.2 is important. If $c$ is specified by $c(x,y,x^{\prime},y^{\prime})=(x-x^{\prime})(y-y^{\prime})$ (or equivalently, if the objective is $\mathrm{Cov}(X,Y)$; c.f. Example 3.3), then the minimizer is the antimonotone coupling instead of the diamond transport.
@@ 6.2 A class of type-XX cost functions with convex QOT
for $\alpha\in(0,1/2]$. Specifically, one of our assumptions in this result is that the quadratic-form transport cost is convex in the transport plan $\pi$. Schoenberg’s theory of complete monotonicity provides a convenient sufficient condition for such convexity. A nonnegative continuous function $\phi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is completely monotone if $\phi$ is $C^{\infty}$ on $(0,\infty)$ and satisfies $(-1)^{n}\phi^{(n)}(u)\geqslant 0$ for $n\geqslant 0,\,u>0$ (see (Berg et al., 1984, Section 4.6)). In particular, $\phi$ is bounded and decreasing. We use the standard calculus notation $\phi^{\prime}$ for the first derivative $\phi^{(1)}$, which should not be confused with the apostrophe in $x^{\prime},y^{\prime}$.
@@
for some completely monotone function $\phi:\mathbb{R}_{+}\to\mathbb{R}$ satisfying
@@
The cost function (6.4) can be written as $c(x,y,x^{\prime},y^{\prime})=h(|x-x^{\prime}|,|y-y^{\prime}|)$, where $h(u,v)=\phi(u^{2})\phi(v^{2})$. Since $u,v\geqslant 0$ and $\phi^{\prime}\leqslant 0$, we have $h_{uv}\geqslant 0$ and hence h-h is submodular. By arguing in the same way as Remark 6.3, we see that under the setting of Theorem 6.5, if $\mu$ is a location-scale transform of $\nu$, then the comonotone and antimonotone couplings are both maximizers.
@@
    Suppose that $\mu,\nu$ are supported on [1/2,1/2][-1/2,1/2]. Let $\phi(u)=(\beta+u)^{-\gamma}$, where $\gamma>0$ and $\beta>2\gamma+1$. This leads to the cost function
@@
    Suppose that $\mu,\nu$ are supported on [L,L][-L,L] where $L>0$. Let $p\in(1/2,1)$ and $\phi(u)=e^{-\alpha u^{p}}$, where $0<\alpha\leqslant(2p-1)/(2p(4L^{2})^{p})$. The cost function is then given by
@@
We first claim that the resulting QOT problem is convex in $\pi$. Since $\phi$ is completely monotone, the function $\psi(u):=\phi(u^{2})$ is a continuous positive definite function on $\mathbb{R}$ by Schoenberg’s theorem (Theorem 4.6.13 and Example 5.1.3 of Berg et al. (1984)). In this way, we write $c(x,y,x^{\prime},y^{\prime})=\psi(x-x^{\prime})\psi(y-y^{\prime})$ where $\psi$ is positive definite, so $(x,x^{\prime})\mapsto\psi(x-x^{\prime})$ and $(y,y^{\prime})\mapsto\psi(y-y^{\prime})$ are positive definite kernels. By the Schur product theorem, the product of two positive definite kernels is also positive definite (see Theorem 3.1.12 of Berg et al. (1984)). Therefore, $c(x,y,x^{\prime},y^{\prime})$ is a positive definite kernel in the two variables $(x,y),(x^{\prime},y^{\prime})\in\mathbb{R}^{2}$. Proposition 4.1 then implies the QOT problem is convex in $\pi$.
Since the cost function (6.4) is translation-invariant, we may assume that $\mu,\nu$ are both symmetric along 0. We next verify
@@
Consider the classic OT problem with cost function $\tilde{c}$ given by (6.7). Denote a minimizer by $\pi_{\tilde{c}}$. By linearity of the classic OT problem and (6.8), another minimizer is given by the symmetrized version $\hat{\pi}_{\tilde{c}}$ of $\pi_{\tilde{c}}$: the law of the uniform mixture of (X,Y),(X,Y),(X,Y),(X,Y),\,(-X,Y),\,(X,-Y), and (X,Y)(-X,-Y) where $(X,Y)\buildrel\mathrm{law}\over{\sim}\pi_{\tilde{c}}$. Therefore, $\hat{\pi}_{\tilde{c}}$ is symmetric along the $x$ and $y$ axes. By cyclical monotonicity and Lemma 6.8, any minimizer is antimonotone on the first and third quadrants and comonotone on the second and fourth quadrants. This implies that $\hat{\pi}_{\tilde{c}}=\pi_{\mathrm{dia}}$, and hence $\pi_{\mathrm{dia}}$} is a minimizer of the OT problem with cost function $\tilde{c}$.
@@
The final claim on uniqueness follows from the fact that if $\phi$ is non-constant, the cost function $c$ is a strictly positive definite kernel in the two variables $(x,y),(x^{\prime},y^{\prime})\in\mathbb{R}^{2}$ (Theorem 3’ of Schoenberg (1938)), and hence the transport cost is strictly convex in $\pi$ with a unique minimizer. ∎
@@ 6.3 The $q$-rectangular cost function
Fix $q\in(1,2]$ and assume without loss of generality that $\mu,\nu$ are both symmetric around 0. Assume first that $\mu,\nu$ are supported in [L,L][-L,L] for some $L>0$. For $\alpha>0$, consider the cost function
@@
uniformly on $L,L]^{4}$ as $\alpha\to 0$. Therefore, the QOT problem with cost function $(|x-x^{\prime}|^{q}+|y-y^{\prime}|^{q})^{2}/2$ also has $\pi_{\mathrm{dia}}$ as a minimizer. Removing the QOT-irrelevant terms, we see that this is also the case for the cost function $|(x-x^{\prime})(y-y^{\prime})|^{q}$, as desired.